# Semantic Structure-Aware Generative Attacks for Enhanced Adversarial Transferability

### v3

@@ 4.2 Evaluation settings
where Top-1 Accuracy [86, 82, 83] measures the overall proportion of correctly classified samples under clean or adversarial conditions. It serves as a global performance indicator to assess degraded performance after the attack, orthogonal to FR, ASR, and ACR. Attack Success Rate (ASR) [51, 50] is a subset of FR, which measures the proportion of originally correctly classified samples that are turned into misclassified ones by the adversarial attack. It directly reflects targeted misclassification. Fooling Rate (FR) [29, 48] quantifies the proportion of adversarial examples that cause a change in the model’s prediction, regardless of correctness. It reflects how often the attack disrupts the original decision and is used as transferability. Accidental Correction Rate (ACR), also a subset of FR, is a novel metric that quantifies how often misclassified samples are “accidentally” corrected by adversarial perturbations. This unintended side effect provides insight into model uncertainty and boundary behavior. For SS and OD tasks, we use the standard mIoU and mAP50 metrics, respectively.
Figure 4: Qualitative results. Our semantic structure-aware attack successfully guides the generator to focus perturbations particularly on the semantically salient regions, effectively fooling the victim classifier. Left: (a) benign input image, (b) generated perturbation (normalized for visual purposes only), (c) unbounded adversarial image, and (d) bounded adversarial image across CUB-200-2011 [67], Stanford Cars [33], FGVC Aircraft [45], and ImageNet-1K [55] domains. The label on top (green) and bottom (orange) denotes the correct label and prediction after the attack, respectively. Center: We highlight that our method induces Grad-CAM [56] to focus on drastically different regions in our adversarial examples compared to both the benign image and the adversarial examples crafted by the baseline [86]. Moreover, our approach noticeably spreads and reduces the high activation regions observed in the benign and baseline cases, enhancing the transferability of our adversarial perturbations. Right: Cross-task prediction results (SS on top, OD on bottom). Our approach further disrupts the victim models by triggering higher false positive rates and wrong class label predictions. Please see the Appendix for more.
@@
Method Cross-setting (Avg.) ϵtest=subscriptitalic-$\epsilon_{\textrm{test}}=$ 2 4 6 8 10 16
Baseline [86] Domain (Cls.; %) 90.15 87.35 78.73 64.95 51.07 24.87
@@ 4.3 Experimental Results
Table S5 reports cross-model evaluations across diverse architectures. As shown, augmenting each baseline generative attack with our method yields consistent improvements in attack success. While these results confirm the orthogonality and efficacy of our framework, we observe that CLIP-based approaches with objectives akin to ours, e.g. PDCL-Attack [83], yield only marginal improvements when combined with our method. We conjecture that optimizing for divergence in CLIP’s high-dimensional semantic embedding space may override or dilute the local structural consistency enforced by our early-block feature distillation, thereby attenuating the incremental gains from preserving fine-grained object contours and textures. Table 3 presents cross-domain transfer results. Incorporating our Mean Teacher smoothing and early-block distillation consistently enhances performance across unseen domains and architectures, demonstrating the broad applicability of our framework beyond the source data distribution. Table 3 evaluates transfer to SS and OD tasks, which were not seen during training. Even in these settings, our feature-distillation loss applied to the Mean Teacher generator produces measurable gains. We attribute this cross-task generalization to our label-agnostic training pipeline and further validate that our method can be integrated with alternative generator architectures beyond ResNet. Qualitative results in Fig. 4 further confirm that our method produces more pronounced adversarial effects on target objects. See details on U-Net in the Appendix.
@@ Appendix A Additional Related Work
For years, iterative gradient-based attacks have become a cornerstone of adversarial research. Methods such as Projected Gradient Descent (PGD) [44] extend the Fast Gradient Sign Method by applying multiple small, $\ell_{\infty}$-bounded steps; Momentum Iterative FGSM (MI-FGSM) [15] further stabilizes updates via accumulated momentum; Diverse Input FGSM (DI-FGSM) [76] injects random resizing and padding at each iteration; and Translation-Invariant FGSM (TI-FGSM) [17] averages gradients over shifted inputs to enhance spatial robustness. More advanced variants even incorporate feature-space objectives to target intermediate representations [85].
@@
In parallel, more efficient generative model-based attacks train a feed-forward image-to-image transformation network to synthesize perturbations in a single pass: Universal Adversarial Perturbations (UAP) [47] learn a single image-agnostic noise vector, Generative Adversarial Perturbations (GAP) [51] use a GAN framework to produce highly transferable noise maps (added to the input images), and AdvGAN [74] leverages GANs for image-dependent attacks that balance stealth and speed. Together, these two paradigms offer complementary trade-offs between precision, transferability, and inference efficiency.
In this vein, recent generative model-based untargeted attack methods [50, 29, 86, 1, 82, 83, 48] have further added techniques to enhance the transferability of the crafted adversarial examples by incorporating surrogate model’s output logit-level and mid-level feature-level separation, frequency domain manipulation, vision-language model guidance, and heuristic selection of one effective neuron-level generator among a pool of multiple generators. However, none of these works have dealt with directly manipulating the generative feature space to improve the transferability of AEs. To address this, we uncover the correlation between generative features and adversarial transferability of the output AEs in this work.
@@
Along with ResNet [24], U-Net [54] is another effective network architecture comprising a symmetric encoder–decoder with skip connections, fusing low- and high-level features to preserve fine-grained details, which are ideal when perturbations must tightly follow object boundaries. By contrast, a ResNet generator stacks residual blocks with identity shortcuts, thus building deep hierarchical representations that emphasize global context. Although U-Net decoders add computational overhead, they can produce sharper, pixel-accurate noise, while ResNet backbones scale more efficiently and excel at generating broadly distributed perturbations. Ultimately, the choice of generator architecture hinges on the desired trade-off between pixel-level fidelity, attack transferability, and inference speed.
In the context of generative adversarial attack, GAP [51] first demonstrated that U-Net can serve as a perturbation generator with a lower inference time cost than that with ResNet. However, the authors of [51] also stated that ResNet, in general, outperforms U-Net in attack transferability. In this work, we demonstrate that our method of anchoring perturbation generation on early-intermediate features can also be applied to a different generator architecture than ResNet, namely U-Net [54].
Specifically, we note the differences between the U-Net and ResNet generators in detail. Due to the symmetric encoder-decoder design in the U-Net, there is only one feature block at the bottleneck that we can employ as the intermediate block feature, as opposed to six in ResNet. Given this architecture, we applied our distillation on this bottleneck feature for the cross-task experiments in Tables 3 and 6, where we observe improvements in attack transferability across domains, models, and tasks with the addition of our components. While we remark noticeable improvements over the ResNet in cross-task settings, ResNet still remains at least comparable or superior in cross-domain and model settings. Given this observation, we remark that while U-Net can still be leveraged as a generative model for transfer-based attacks, further research on boosting U-Net-based attack transferability may be necessary. We also note that proper U-Net [54] reference will be added to the manuscript.
@@ Appendix B Our method distinction from previous works
GAP [51] - - - - -
CDA [70] - - - - -
LTP [29] - - - - -
BIA [86] - - - -
GAMA [1] - - - -
FACL [82] - - -
PDCL [83] - - - -
@@
Previous works have targeted various stages of the generative attack pipeline [51, 50, 29, 86], including input data augmentation, pixel-level perturbation, surrogate model’s logit- and feature-level manipulations. Nonetheless, no work has yet explicitly manipulated the internal features of the generative model to enhance transferability. In this work, we investigate how internal feature representations within generative models can be harnessed to enhance the transferability of AEs.
@@
In the figure below (Fig. S2), we directly compare the feature activation maps and the added adversarial noise per block (absolute difference of the input and output of each block) of Ours against the baseline [86]. We particularly focus on the intermediate residual blocks (“Residual Learning"), as most of the adversarial noise is generated in these blocks [86], and the preceding (“Downsampling" layers) and succeeding blocks (“Upsampling" layers) serve to simply adjust the spatial resolution of the feature maps.
@@
Figure S2: Comparison of the input image (column 1), the feature activation maps, and the added adversarial noise between two residual blocks (absolute difference between the feature maps of two residual blocks) of Ours (columns 4, 5) against the baseline [86] (columns 2, 3). Although the baseline’s feature maps more vividly emphasize object boundaries and contours, this focus actually prevents perturbations from appearing in those highlighted regions. By contrast, our method produces relatively less pronounced early features than those of the baseline, yet focuses more perturbation power on object-salient regions towards the later blocks than the baseline, thereby allowing adversarial noise to be dispersed directly on and around those salient regions, as opposed to the baseline. For the feature activation maps and the added noise, the brighter and darker (respectively), the higher the value.
@@
We further summarize our findings and approach as follows. We empirically discover that existing generative works preserve the semantic integrity of the benign input image at the early intermediate blocks better than the later blocks, under the assumption that the downsampling blocks merely serve as feature extractors [86]. To better structure adversarial noise generation in the intermediate blocks, we formulate our method to maintain at least coarse semantic structures in the early intermediate blocks, thereby yielding much more perturbed features towards the end of the intermediate stage, which then results in enhanced transferability of the crafted AEs compared to the baseline. Through extensive cross-setting evaluations, we validate our approach of tuning the progression of adversarial noise generation in the generator’s feature level as a compatible method to the existing generative attack framework [50, 29, 86, 1, 82, 83, 48] without much overheads.
@@ C.1 Details on the evaluation settings
We define three black-box evaluation scenarios that differ in the attacker’s knowledge of target data and models. In the cross-model setting, the adversary has access to the same data distribution used to train the unseen target models but must craft attacks using a substitute model rather than querying the targets directly. In the cross-domain setting, the attacker works solely with out-of-domain data and has no access to target-domain datasets (e.g., CUB-200-2011 [67], Stanford Cars [33], FGVC Aircraft [45]) or the ability to query target-domain models such as ResNet-50 [24], SE-Net, or SE-ResNet101 [25]. Finally, in the cross-task setting, likewise, the adversary is completely agnostic to the target’s data, models, and even the task itself, representing the strict black-box challenge.
@@
we selected a total of 22 different model architectures that span from CNNs to Vision Mamba variants, whose pre-trained model weights are available openly through TorchVision [46], Timm [72], and the proprietary GitHub repositories. We list the sources in Table S2.
@@
          Image Classification          ResNet50 [24]          TorchVision [46]
         ResNet152 [24]          TorchVision [46]
         Dense121 [26]          TorchVision [46]
         Dense169 [26]          TorchVision [46]
         InceptionV3 [59]          TorchVision [46]
         RegNetY [53]          TorchVision [46]
         MNAsNet [61]          TorchVision [46]
         SqueezeNet [28]          TorchVision [46]
         EfficientV2 [60]          TorchVision [46]
         HRNet_W32 [69]          Timm [72]
         ConvNeXt [42]          TorchVision [46]
         ResNeXt [79]          TorchVision [46]
         ViT-B/16 [18]          TorchVision [46]
         ViT-L/16 [18]          TorchVision [46]
         Swin-B/16 [41]          TorchVision [46]
         DeiT-B [64]          Timm [72]
         BEiT-B [4]          Timm [72]
         EfficientViT [6]          Timm [72]
         MLP-Mixer-B [63]          Timm [72]
         ConvMixer-B [65]          Timm [72]
         Vision Mamba-B [89]          https://github.com/hustvl/Vim
         MambaVision-B [23]          https://github.com/NVlabs/MambaVision
          SemSeg          DeepLabV3+ [11]          https://github.com/VainF/DeepLabV3Plus-Pytorch
         SegFormer [78]          https://github.com/NVlabs/SegFormer
          ObjDet          Faster R-CNN [20]          https://github.com/facebookresearch/detectron2
         DETR [8]          https://github.com/facebookresearch/detr

### v4

@@ 4.2 Evaluation settings
where Top-1 Accuracy Zhang et al. (2022); Yang et al. (2024a, b) measures the overall proportion of correctly classified samples under clean or adversarial conditions. It serves as a global performance indicator to assess degraded performance after the attack, orthogonal to FR, ASR, and ACR. Attack Success Rate (ASR) Poursaeed et al. (2018); Naseer et al. (2019) is a subset of FR, which measures the proportion of originally correctly classified samples that are turned into misclassified ones by the adversarial attack. It directly reflects targeted misclassification. Fooling Rate (FR) kanth Nakka and Salzmann (2021); Nakka and Alahi (2025) quantifies the proportion of adversarial examples that cause a change in the model’s prediction, regardless of correctness. It reflects how often the attack disrupts the original decision and is used as transferability. Accidental Correction Rate (ACR), also a subset of FR, is a novel metric that quantifies how often misclassified samples are “accidentally” corrected by adversarial perturbations. This unintended side effect provides insight into model uncertainty and boundary behavior. For SS and OD tasks, we use the standard mIoU and mAP50 metrics, respectively.
Figure 4: Qualitative results. Our semantic structure-aware attack successfully guides the generator to focus perturbations particularly on the semantically salient regions, effectively fooling the victim classifier. Left: (a) benign input image, (b) generated perturbation (normalized for visual purposes only), (c) unbounded adversarial image, and (d) bounded adversarial image across CUB-200-2011 Wah et al. (2011), Stanford Cars Krause et al. (2013), FGVC Aircraft Maji et al. (2013), and ImageNet-1K Russakovsky et al. (2015) domains. The label on top (green) and bottom (orange) denotes the correct label and prediction after the attack, respectively. Center: We highlight that our method induces Grad-CAM Selvaraju et al. (2017) to focus on drastically different regions in our adversarial examples compared to both the benign image and the adversarial examples crafted by the baseline Zhang et al. (2022). Moreover, our approach noticeably spreads and reduces the high activation regions observed in the benign and baseline cases, enhancing the transferability of our adversarial perturbations. Right: Cross-task prediction results ($S$ on top, OD on bottom). Our approach further disrupts the victim models by triggering higher false positive rates and wrong class label predictions. Please see the Appendix for more.
@@
Method Cross-setting (Avg.) $\epsilon_{\textrm{test}}=$ 2 4 6 8 10 16
Baseline Zhang et al. (2022) Domain (Cls.; %) 90.15 87.35 78.73 64.95 51.07 24.87
@@ 4.3 Experimental Results
Table S5 reports cross-model evaluations across diverse architectures. As shown, augmenting each baseline generative attack with our method yields consistent improvements in attack success. While these results confirm the orthogonality and efficacy of our framework, we observe that CLIP-based approaches with objectives akin to ours, e.g. PDCL-Attack Yang et al. (2024b), yield only marginal improvements when combined with our method. We conjecture that optimizing for divergence in CLIP’s high-dimensional semantic embedding space may override or dilute the local structural consistency enforced by our early-block feature distillation, thereby attenuating the incremental gains from preserving fine-grained object contours and textures. Table 3 presents cross-domain transfer results. Incorporating our Mean Teacher smoothing and early-block distillation consistently enhances performance across unseen domains and architectures, demonstrating the broad applicability of our framework beyond the source data distribution. Table 3 evaluates transfer to SS and OD tasks, which were not seen during training. Even in these settings, our feature-distillation loss applied to the Mean Teacher generator produces measurable gains. We attribute this cross-task generalization to our label-agnostic training pipeline and further validate that our method can be integrated with alternative generator architectures beyond ResNet. Qualitative results in Fig. 4 further confirm that our method produces more pronounced adversarial effects on target objects. See details on U-Net in the Appendix.
@@ Appendix A Additional Related Work
For years, iterative gradient-based attacks have become a cornerstone of adversarial research. Methods such as Projected Gradient Descent (PGD) Madry et al. (2017b) extend the Fast Gradient Sign Method by applying multiple small, $\ell_{\infty}$-bounded steps; Momentum Iterative FGSM (MI-FGSM) Dong et al. (2018) further stabilizes updates via accumulated momentum; Diverse Input FGSM (DI-FGSM) Xie et al. (2019a) injects random resizing and padding at each iteration; and Translation-Invariant FGSM (TI-FGSM) Dong et al. (2019b) averages gradients over shifted inputs to enhance spatial robustness. More advanced variants even incorporate feature-space objectives to target intermediate representations Zhang et al. (2021).
@@
In parallel, more efficient generative model-based attacks train a feed-forward image-to-image transformation network to synthesize perturbations in a single pass: Universal Adversarial Perturbations (UAP) Moosavi-Dezfooli et al. (2017) learn a single image-agnostic noise vector, Generative Adversarial Perturbations (GAP) Poursaeed et al. (2018) use a GAN framework to produce highly transferable noise maps (added to the input images), and AdvGAN Xiao et al. (2018) leverages GANs for image-dependent attacks that balance stealth and speed. Together, these two paradigms offer complementary trade-offs between precision, transferability, and inference efficiency.
In this vein, recent generative model-based untargeted attack methods Naseer et al. (2019); kanth Nakka and Salzmann (2021); Zhang et al. (2022); Aich et al. (2022); Yang et al. (2024a, b); Nakka and Alahi (2025) have further added techniques to enhance the transferability of the crafted adversarial examples by incorporating surrogate model’s output logit-level and mid-level feature-level separation, frequency domain manipulation, vision-language model guidance, and heuristic selection of one effective neuron-level generator among a pool of multiple generators. However, none of these works have dealt with directly manipulating the generative feature space to improve the transferability of AEs. To address this, we uncover the correlation between generative features and adversarial transferability of the output AEs in this work.
@@
Along with ResNet He et al. (2016), U-Net Ronneberger et al. (2015) is another effective network architecture comprising a symmetric encoder–decoder with skip connections, fusing low- and high-level features to preserve fine-grained details, which are ideal when perturbations must tightly follow object boundaries. By contrast, a ResNet generator stacks residual blocks with identity shortcuts, thus building deep hierarchical representations that emphasize global context. Although U-Net decoders add computational overhead, they can produce sharper, pixel-accurate noise, while ResNet backbones scale more efficiently and excel at generating broadly distributed perturbations. Ultimately, the choice of generator architecture hinges on the desired trade-off between pixel-level fidelity, attack transferability, and inference speed.
In the context of generative adversarial attack, GAP Poursaeed et al. (2018) first demonstrated that U-Net can serve as a perturbation generator with a lower inference time cost than that with ResNet. However, the authors of Poursaeed et al. (2018) also stated that ResNet, in general, outperforms U-Net in attack transferability. In this work, we demonstrate that our method of anchoring perturbation generation on early-intermediate features can also be applied to a different generator architecture than ResNet, namely U-Net Ronneberger et al. (2015).
Specifically, we note the differences between the U-Net and ResNet generators in detail. Due to the symmetric encoder-decoder design in the U-Net, there is only one feature block at the bottleneck that we can employ as the intermediate block feature, as opposed to six in ResNet. Given this architecture, we applied our distillation on this bottleneck feature for the cross-task experiments in Tables 3 and 6, where we observe improvements in attack transferability across domains, models, and tasks with the addition of our components. While we remark noticeable improvements over the ResNet in cross-task settings, ResNet still remains at least comparable or superior in cross-domain and model settings. Given this observation, we remark that while U-Net can still be leveraged as a generative model for transfer-based attacks, further research on boosting U-Net-based attack transferability may be necessary. We also note that proper U-Net Ronneberger et al. (2015) reference will be added to the manuscript.
@@ Appendix B Our method distinction from previous works
GAP Poursaeed et al. (2018) - - - - -
CDA Wang et al. (2022) - - - - -
LTP kanth Nakka and Salzmann (2021) - - - - -
BIA Zhang et al. (2022) - - - -
GAMA Aich et al. (2022) - - - -
FACL Yang et al. (2024a) - - -
PDCL Yang et al. (2024b) - - - -
@@
Previous works have targeted various stages of the generative attack pipeline Poursaeed et al. (2018); Naseer et al. (2019); kanth Nakka and Salzmann (2021); Zhang et al. (2022), including input data augmentation, pixel-level perturbation, surrogate model’s logit- and feature-level manipulations. Nonetheless, no work has yet explicitly manipulated the internal features of the generative model to enhance transferability. In this work, we investigate how internal feature representations within generative models can be harnessed to enhance the transferability of AEs.
@@
In the figure below (Fig. S2), we directly compare the feature activation maps and the added adversarial noise per block (absolute difference of the input and output of each block) of Ours against the baseline Zhang et al. (2022). We particularly focus on the intermediate residual blocks (“Residual Learning"), as most of the adversarial noise is generated in these blocks Zhang et al. (2022), and the preceding (“Downsampling" layers) and succeeding blocks (“Upsampling" layers) serve to simply adjust the spatial resolution of the feature maps.
@@
Figure S2: Comparison of the input image (column 1), the feature activation maps, and the added adversarial noise between two residual blocks (absolute difference between the feature maps of two residual blocks) of Ours (columns 4, 5) against the baseline Zhang et al. (2022) (columns 2, 3). Although the baseline’s feature maps more vividly emphasize object boundaries and contours, this focus actually prevents perturbations from appearing in those highlighted regions. By contrast, our method produces relatively less pronounced early features than those of the baseline, yet focuses more perturbation power on object-salient regions towards the later blocks than the baseline, thereby allowing adversarial noise to be dispersed directly on and around those salient regions, as opposed to the baseline. For the feature activation maps and the added noise, the brighter and darker (respectively), the higher the value.
@@
We further summarize our findings and approach as follows. We empirically discover that existing generative works preserve the semantic integrity of the benign input image at the early intermediate blocks better than the later blocks, under the assumption that the downsampling blocks merely serve as feature extractors Zhang et al. (2022). To better structure adversarial noise generation in the intermediate blocks, we formulate our method to maintain at least coarse semantic structures in the early intermediate blocks, thereby yielding much more perturbed features towards the end of the intermediate stage, which then results in enhanced transferability of the crafted AEs compared to the baseline. Through extensive cross-setting evaluations, we validate our approach of tuning the progression of adversarial noise generation in the generator’s feature level as a compatible method to the existing generative attack framework Naseer et al. (2019); kanth Nakka and Salzmann (2021); Zhang et al. (2022); Aich et al. (2022); Yang et al. (2024a, b); Nakka and Alahi (2025) without much overheads.
@@ C.1 Details on the evaluation settings
We define three black-box evaluation scenarios that differ in the attacker’s knowledge of target data and models. In the cross-model setting, the adversary has access to the same data distribution used to train the unseen target models but must craft attacks using a substitute model rather than querying the targets directly. In the cross-domain setting, the attacker works solely with out-of-domain data and has no access to target-domain datasets (e.g., CUB-200-2011 Wah et al. (2011), Stanford Cars Krause et al. (2013), FGVC Aircraft Maji et al. (2013)) or the ability to query target-domain models such as ResNet-50 He et al. (2016), SE-Net, or SE-ResNet101 Hu et al. (2018). Finally, in the cross-task setting, likewise, the adversary is completely agnostic to the target’s data, models, and even the task itself, representing the strict black-box challenge.
@@
we selected a total of 22 different model architectures that span from CNNs to Vision Mamba variants, whose pre-trained model weights are available openly through TorchVision Marcel and Rodriguez (2010), Timm Wightman , and the proprietary GitHub repositories. We list the sources in Table S2.
@@
Image Classification
   ResNet50 He et al. (2016)    TorchVision Marcel and Rodriguez (2010)
   ResNet152 He et al. (2016)    TorchVision Marcel and Rodriguez (2010)
   Dense121 Huang et al. (2017)    TorchVision Marcel and Rodriguez (2010)
   Dense169 Huang et al. (2017)    TorchVision Marcel and Rodriguez (2010)
   InceptionV3 Szegedy et al. (2016)    TorchVision Marcel and Rodriguez (2010)
   RegNetY Radosavovic et al. (2020)    TorchVision Marcel and Rodriguez (2010)
   MNAsNet Tan et al. (2019)    TorchVision Marcel and Rodriguez (2010)
   SqueezeNet Iandola et al. (2016)    TorchVision Marcel and Rodriguez (2010)
   EfficientV2 Tan and Le    TorchVision Marcel and Rodriguez (2010)
   HRNet_W32 Wang et al. (2020)    Timm Wightman
   ConvNeXt Liu et al. (2022)    TorchVision Marcel and Rodriguez (2010)
   ResNeXt Xie et al. (2017)    TorchVision Marcel and Rodriguez (2010)
   ViT-B/16 Dosovitskiy et al. (2021)    TorchVision Marcel and Rodriguez (2010)
   ViT-L/16 Dosovitskiy et al. (2021)    TorchVision Marcel and Rodriguez (2010)
   Swin-B/16 Liu et al. (2021)    TorchVision Marcel and Rodriguez (2010)
   DeiT-B Touvron et al. (2021)    Timm Wightman
   BEiT-B Bao et al. (2021)    Timm Wightman
   EfficientViT Cai et al. (2022)    Timm Wightman
   MLP-Mixer-B Tolstikhin et al. (2021)    Timm Wightman
   ConvMixer-B Trockman and Kolter (2022)    Timm Wightman
   Vision Mamba-B Zhu et al. (2024)    https://github.com/hustvl/Vim
   MambaVision-B Hatamizadeh and Kautz (2025)    https://github.com/NVlabs/MambaVision
SemSeg
   DeepLabV3+ Chen et al. (2018)    https://github.com/VainF/DeepLabV3Plus-Pytorch
   SegFormer Xie et al. (2021)    https://github.com/NVlabs/SegFormer
ObjDet
   Faster R-CNN Girshick (2015)    https://github.com/facebookresearch/detectron2
   DETR Carion et al. (2020)    https://github.com/facebookresearch/detr